The smallest angle in a triangle is half the size of the largest. This middle angle is 25 degrees less than the largest. Find the measures of the three angles.


The hand soap, Dove, is 1/4 moisturizing cream. How much of a 4.5 oz bar is something other than moisturizing cream?

Answer by longjonsilver(2297)

Let smallest angle = x
so largest angle = 2x
Middle angle is 2x-25

Being a triangle, all 3 add upto 180, so

x + 2x + (2x-25) = 180
5x - 25 = 180
5x = 205
x = 205/5
x = 41

So, smallest angle = 41
largest angle = 82
middle angle = 57

check... 41+82+57 is 180

jon.